# QuantumBind-RBFE: Accurate Relative Binding Free Energy Calculations Using Neural Network Potentials

### v1

@@ 3.2 Identifying the top compounds in a series
Results in (Figure 5) indicate that AceForce 1.0 generally performs as well as or better than the other methods in identifying the most potent compounds. For example, in datasets such as BACE, CDK2, P38, and TYK2, AceForce 1.0 successfully identified at least 50% of the top 30% compounds (59%, 52%, 50%, and 84%, respectively), outperforming GAFF2 runs in all cases. However, the identification of top compounds by AceForce 1.0 is less effective in some datasets, such as JNK1, MCL1, and THROMBIN, where the top 30% identification is considerably lower. We believe that this metric is particularly valuable for larger datasets such as BACE, MCL1, and P38. For these datasets, AceForce 1.0 generally shows improved performance compared to GAFF2, though with variability across targets: BACE and P38 exhibit clear improvements, while MCL1 remains a challenge.
In terms of top 5 identification, the differences between methods were generally small. For MCL1, AceForce 1.0 placed one ligand in the top 5, while GAFF2 and OPLS4 identified two and three, respectively. In most other cases, the same number of top-ranked molecules were identified, with discrepancies of no more than a single ligand when comparing AceForce 1.0 to OPLS4 (Figure 5). Notably, even in cases where AceForce 1.0 exhibited higher overall error rates compared to OPLS4, its compound prioritization remained comparable.
Figure 5: Identification of top compounds across datasets. The left plot shows the accuracy in identifying compounds within the top 30%, while the right plot focuses on the top 5 compounds in each series. Results are compared across the evaluated methods: AceForce 1.0, GAFF2, and OPLS4 with FEP+.
@@ 3.3 Timesteps: 1fs vs 2fs
With the AceForce 1.0 model for TorchMD-Net, we observed improved stability at a 2fs timestep, significantly enhancing the computational speed of NNP/MM simulations. To evaluate whether this larger timestep maintains accuracy, we conducted RBFE calculations on a subset of the JACS dataset at a timestep of both 1 and 2fs. Our findings (Figure 6 and Supporting TablesS10 and S11) indicate that 2 fs timestep simulations demonstrate comparable accuracy to 1 fs runs for most targets. In these cases, RMSE and Kendall tau correlation values remained nearly identical. The impact on top compound identification is shown in Supporting Table S12. Considering the variability between runs, we can conclude that it is possible to run at simulations at 2 fs.
Figure 6: Comparison of AceForce 1.0 model accuracy at 1fs and 2fs timesteps. The top panel shows Root Mean Squared Error (RMSE) and the bottom panel shows Kendall tau correlation for $\Delta G$ values across a subset of systems. While runs at both timesteps yield similar performance, the 1fs timestep calculations maintain slightly higher accuracy.
@@ 3.4 Comparison with ANI-2x
We compare now the performance of AceForce to ANI-2x in terms of accuracy.22 From the JACS dataset, we computed several edges for CDK2, JNK1, P38, and TYK2. Due to limitations in ANI-2x’s atom type and charge handling, the datasets for JNK1 and CDK2 are not complete, so we compare only the same edges that both methods could calculate. In contrast, the P38 and TYK2 datasets are complete for both models, allowing for a more direct assessment of their performance. For this reason, a comparison between $\Delta G$ values is only done for P38 and TYK2. RMSE and MAE values for the computed $\Delta\Delta G$ edges are presented in Figure 7. Additionally, table S9 presents a summary of the $\Delta G$ results, while Table S8 provides the $\Delta\Delta G$ statistics for all four targets. For $\Delta\Delta G$ calculations on P38, AceForce achieves a slightly lower RMSE of 1.04 kcal/mol compared to ANI-2x’s 1.17 kcal/mol, alongside a lower MAE (0.84 vs. 0.91 kcal/mol). Additionally, AceForce shows a higher Kendall tau correlation (0.69 vs. 0.59). On TYK2, both models deliver similar accuracy, with AceForce showing a slightly lower RMSE (0.55 vs. 0.56 kcal/mol) and a modestly higher Kendall tau (0.73 vs. 0.67). For CDK2, AceForce yields an RMSE of 1.03 kcal/mol, slightly higher than ANI-2x’s 0.83 kcal/mol, and a Kendall tau of 0.46 compared to ANI-2x’s 0.62, indicating that ANI-2x captures the ranking of these particular ligands more effectively. Conversely, on JNK1, AceForce attains an RMSE of 0.92 kcal/mol vs. ANI-2x’s 0.90 kcal/mol, and a higher Kendall tau (0.46 vs. 0.43), suggesting a marginal ranking advantage for AceForce on this target. For absolute binding free energies ($\Delta G$), AceForce again demonstrates improved overall performance. On P38, AceForce achieves an RMSE of 0.77 kcal/mol and a Kendall tau of 0.72, and ANI-2x (RMSE 0.93, Kendall tau 0.58). Similarly, on TYK2, AceForce reaches an RMSE of 0.47 kcal/mol and Kendall tau of 0.85, slightly surpassing ANI-2x’s RMSE of 0.50 and Kendall tau of 0.82. Overall, these findings indicate that AceForce generally perform better across the evaluated systems with the additional benefit that it is not limited to those.
Figure 7: Comparison of AceForce and ANI-2x calculations for $\Delta\Delta G$ values.
@@ 3.5 Going beyond the limits of the model
PTP1B is a protein target included in the JACS dataset, characterized by ligands uniformly carrying a -2 charge. AceForce 1.0, however, has been trained on molecules with charges limited to -1, 0, and +1, making its applicability for this target troublesome. This limitation of AceForce will be solved in version 1.1. We conducted RBFE calculations for PTP1B to assess the model’s performance under such constraints. As anticipated, the predictions exhibited significant errors, with RMSE exceeding 3 kcal/mol for both $\Delta\Delta G$ and $\Delta G$ values and two out of the 49 runs crashing. Furthermore, the correlation performance was notably poor, reflected by a Kendall tau coefficient of -0.31, and the method failed to correctly identify any of the top ligands in the dataset (Figure8). This outcome aligns with expectations, emphasizing the need for further training of the model on an extended range of ligand charges to improve its generalization and predictive capabilities.
Figure 8: Top: Example on one of the ligands of the PTP1B dataset, highlighting the -1 charge at each of the two COO- groups. Bottom: scatterplot for the relative and absolute binding free energy calculations. Relevant metrics are reported on top of each plot
@@ 4 Conclusions
We are presenting our NNP/MM approach to perform RBFE calculations. We are using AceForce 1.0, a novel NNP model that supports a broader range of atom elements and charged molecules. To assess its accuracy, we benchmarked AceForce 1.0 using the JACS dataset27 (without PTP1B due to charge limitations) and compared it to an MM-based approach as well as the industry state-of-the-art alternative in OPLS4 with FEP+.28 AceForce 1.0 demonstrated overall improved performance to MM-based calculations like GAFF2. These improvements highlight NNP potentials as an accurate alternative to MM approaches for RBFE calculations. In comparison to OPLS4, the state-of-the-art approach in the industry, AceForce 1.0 provided competitive or superior correlation performance on half of the evaluated systems but underperformed on the rest. These results indicate that further training of the potential may be needed to consistently achieve SOTA accuracy across all targets. Additionally, when compared to ANI-2x, the initial NNP tested in our previous work,22 AceForce 1.0 delivered better accuracy overall.
@@
Future work will focus on testing and developing new versions of AceForce with an expanded training dataset to enhance predictive accuracy and general applicability. We also aim to enable stable simulations at larger timesteps, potentially 3fs or even 4fs, to reduce the computational cost of NNP/MM calculations to levels comparable with MM-based methods. Additionally, we plan to expand RBFE studies to include other benchmark datasets.
@@ 5 Acknowledgement
The authors thank the volunteers of GPUGRID.net for donating computing time. This project has received funding from the Torres-Quevedo Programme from the Spanish National Agency for Research (PTQ2020-011145 / AEI / 10.13039/501100011033).
@@ 6 Conflict of interests
7 Code and Model availability
AceForce models are available from HuggingFace https://huggingface.co/Acellera/AceForce .
Examples of how to use it are available: 1) Run ML potential molecular simulations of a small molecule using ACEMD https://software.acellera.com/acemd/nnp.html. 2) For a tutorial on running mixed protein-ligand simulations, refer to NNP/MM https://software.acellera.com/acemd/nnpmm.html.
A tutorial on how to run the RBFE calculations, input data and structures are available at https://github.com/Acellera/quantumbind_rbfe.

### v2

@@ 3.2 Identifying the top compounds in a series
Results (Figure 5 and Table S8) indicate that AceFF 1.0 generally performs as well as or better than the other methods in identifying the most potent compounds. For example, in datasets such as BACE, CDK2, P38, and TYK2, AceFF 1.0 successfully identified at least 50% of the top 30% compounds (59%, 52%, 50%, and 84%, respectively), outperforming GAFF2 runs in all cases. However, the identification of top compounds by AceFF 1.0 is less effective in some datasets, such as JNK1, MCL1, and THROMBIN, where the top 30% identification is considerably lower. We believe that this metric is particularly valuable for larger datasets such as BACE, MCL1, and P38. For these datasets, AceFF 1.0 generally shows improved performance compared to GAFF2, though with variability across targets: BACE and P38 exhibit clear improvements, while MCL1 remains a challenge.
In terms of top 5 identification, the differences between methods were generally small. For MCL1, AceFF 1.0 placed one ligand in the top 5, while GAFF2 and OPLS4 identified two and three, respectively. In most other cases, the same number of top-ranked molecules were identified, with discrepancies of no more than a single ligand when comparing AceFF 1.0 to OPLS4 (Figure 5). Notably, even in cases where AceFF 1.0 exhibited higher overall error rates compared to OPLS4, its compound prioritization remained comparable.
Figure 5: Identification of top compounds across datasets. The left plot shows the accuracy in identifying compounds within the top 30%, while the right plot focuses on the top 5 compounds in each series. Results are compared across the evaluated methods: AceFF 1.0, GAFF2, and OPLS4 with FEP+.
@@ 3.3 Timesteps: 1fs vs 2fs
With the AceFF 1.0 model for TorchMD-Net, we observed improved stability at a 2fs timestep, significantly enhancing the computational speed of NNP/MM simulations.(Figure S2) To evaluate whether this larger timestep maintains accuracy, we conducted RBFE calculations on a subset of the JACS dataset at a timestep of both 1 and 2fs. Our findings (Figure 6 and Tables S9 and S10) indicate that 2 fs timestep simulations demonstrate comparable accuracy to 1 fs runs for most targets. In these cases, RMSE and Kendall tau correlation values remained nearly identical. The impact on top compound identification is shown in Table S11. Considering the variability between runs, we can conclude that it is possible to run at simulations at 2 fs.
Figure 6: Comparison of AceFF 1.0 model accuracy at 1fs and 2fs timesteps. The top panel shows Root Mean Squared Error (RMSE) and the bottom panel shows Kendall tau correlation for $\Delta G$ values across a subset of systems. While runs at both timesteps yield similar performance, the 1fs timestep calculations maintain slightly higher accuracy.
@@ 3.4 Comparison with ANI-2x
We compare now the performance of AceFF to ANI-2x in terms of accuracy.23 From the JACS dataset, we computed several edges for CDK2, JNK1, P38, and TYK2. Due to limitations in ANI-2x’s atom type and charge handling, the datasets for JNK1 and CDK2 are not complete, so we compare only the same edges that both methods could calculate. In contrast, the P38 and TYK2 datasets are complete for both models, allowing for a more direct assessment of their performance. For this reason, a comparison between $\Delta G$ values is only done for P38 and TYK2. RMSE and MAE values for the computed $\Delta\Delta G$ edges are presented in Figure 7. Additionally, Table S12 presents a summary of the $\Delta G$ results, while Table S13 provides the $\Delta\Delta G$ statistics for all four targets. For $\Delta\Delta G$ calculations on P38, AceFF achieves a slightly lower RMSE of 1.04 kcal/mol compared to ANI-2x’s 1.17 kcal/mol, alongside a lower MAE (0.84 vs. 0.91 kcal/mol). Additionally, AceFF shows a higher Kendall tau correlation (0.69 vs. 0.59). On TYK2, both models deliver similar accuracy, with AceFF showing a slightly lower RMSE (0.55 vs. 0.56 kcal/mol) and a modestly higher Kendall tau (0.73 vs. 0.67). For CDK2, AceFF yields an RMSE of 1.03 kcal/mol, slightly higher than ANI-2x’s 0.83 kcal/mol, and a Kendall tau of 0.46 compared to ANI-2x’s 0.62, indicating that ANI-2x captures the ranking of these particular ligands more effectively. Conversely, on JNK1, AceFF attains an RMSE of 0.92 kcal/mol vs. ANI-2x’s 0.90 kcal/mol, and a higher Kendall tau (0.46 vs. 0.43), suggesting a marginal ranking advantage for AceFF on this target. For absolute binding free energies ($\Delta G$), AceFF again demonstrates improved overall performance. On P38, AceFF achieves an RMSE of 0.77 kcal/mol and a Kendall tau of 0.72, and ANI-2x (RMSE 0.93, Kendall tau 0.58). Similarly, on TYK2, AceFF reaches an RMSE of 0.47 kcal/mol and Kendall tau of 0.85, slightly surpassing ANI-2x’s RMSE of 0.50 and Kendall tau of 0.82. Overall, these findings indicate that AceFF generally perform better across the evaluated systems with the additional benefit that it is not limited to those.
Figure 7: Comparison of AceFF and ANI-2x calculations for $\Delta\Delta G$ values.
@@ 3.5 Going beyond the limits of the model
PTP1B is a protein target included in the JACS dataset, characterized by ligands uniformly carrying a -2 charge. AceFF 1.0, however, has been trained on molecules with charges limited to -1, 0, and +1, making its applicability for this target troublesome. This limitation of AceFF will be solved in version 1.1. We conducted RBFE calculations for PTP1B to assess the model’s performance under such constraints. As anticipated, the predictions exhibited significant errors, with RMSE exceeding 3 kcal/mol for both $\Delta\Delta G$ and $\Delta G$ values and two out of the 49 runs crashing. Furthermore, the correlation performance was notably poor, reflected by a Kendall tau coefficient of -0.31, and the method failed to correctly identify any of the top ligands in the dataset (Figure 8). This outcome aligns with expectations, emphasizing the need for further training of the model on an extended range of ligand charges to improve its generalization and predictive capabilities.
Figure 8: Left: Example on one of the ligands of the PTP1B dataset, highlighting the -1 charge at each of the two COO- groups. Right: scatterplot for the relative and absolute binding free energy calculations. Relevant metrics are reported on top of each plot
@@ 4 Conclusions
We are presenting our NNP/MM approach to perform RBFE calculations. We are using AceFF 1.0, a novel NNP model that supports a broader range of atom elements and charged molecules. To assess its accuracy, we benchmarked AceFF 1.0 using the JACS dataset32 (without PTP1B due to charge limitations) and compared it to an MM-based approach as well as the industry state-of-the-art alternative in OPLS4 with FEP+.33 AceFF 1.0 demonstrated overall improved performance to MM-based calculations like GAFF2. These improvements highlight NNP potentials as an accurate alternative to MM approaches for RBFE calculations. In comparison to OPLS4, the state-of-the-art approach in the industry, AceFF 1.0 provided competitive or superior correlation performance on half of the evaluated systems but underperformed on the rest. These results indicate that further training of the potential may be needed to consistently achieve SOTA accuracy across all targets. Additionally, when compared to ANI-2x, the initial NNP tested in our previous work,23 AceFF 1.0 delivered better accuracy overall.
It is important to note, however, that the performance of NNPs is strongly dependent on the diversity of their training datasets. For instance, AceFF 1.0 was trained only on molecules with charges of –1, 0, and +1. Consequently, it does not perform well for systems involving ligands with rare charge states, as observed with the PTP1B dataset. This limitation underscores the need for expanding the training data to encompass a broader chemical space, which we plan to address in future versions of AceFF.
@@
Future work will focus on testing and developing new versions of AceFF with an expanded training dataset to enhance predictive accuracy and general applicability. We also aim to enable stable simulations at larger timesteps, potentially 3fs or even 4fs, to reduce the computational cost of NNP/MM calculations to levels comparable with MM-based methods. Additionally, we plan to expand RBFE studies to include other benchmark datasets.
@@ 5 Acknowledgement
The authors thank the volunteers of GPUGRID.net for donating computing time. This project has received funding from the Torres-Quevedo Programme from the Spanish National Agency for Research (PTQ2023-012967 / AEI / 10.13039/501100011033).
@@ 6 Conflict of interests
7 Author Contributions
F.S.Z wrote the manuscript and performed the RBFE benchmark; S.E.F and S.D built AceFF 1.0 and developed QuantumBind-RBFE; G.D.F designed the project and wrote the manuscript
8 Associated Content
The Supporting Information contains:
  • Alchemical Schedule: A detailed table showing the alchemical schedule of the Softplus Alchemical Potential parameters used for the two legs of the alchemical transformations.
  • Free Energy Statistics: Comprehensive tables summarizing the raw $\Delta G$ and $\Delta\Delta G$ values, along with RMSE, MAE, and Kendall tau correlation metrics for all evaluated targets across the different methods (GAFF2, OPLS4 with FEP+, and AceFF 1.0).
  • Charge Model Comparisons: Detailed comparisons between RESP and AM1BCC charge models, including performance metrics and top identifier analyses.
  • Timestep comparisons: additional tables comparing AceFF at 1 and 2fs timesteps.
  • Comparative Analyses: additional tables comparing AceFF and ANI-2x calculations for both $\Delta G$ and $\Delta\Delta G$ values
  • Simulation speed benchmarks: a figure illustrating the simulation speed for various modes.
9 Data and software availability